# references.bib

### v1

@@ Abstract
During the acquisition of satellite images, there is generally a trade-off between spatial resolution and temporal resolution (acquisition frequency) due to the onboard sensors of satellite imaging systems. High-resolution satellite images are very important for land crop monitoring, urban planning, wildfire management and a variety of applications. It is a significant yet challenging task to achieve high spatial-temporal resolution in satellite imaging. With the advent of diffusion models, we can now learn strong generative priors to generate realistic satellite images with high resolution, which can be utilized to promote the super-resolution task as well. In this work, we propose a novel diffusion-based fusion algorithm called SatDiffMoE that can take an arbitrary number of sequential low-resolution satellite images at the same location as inputs, and fuse them into one high-resolution reconstructed image with more fine details, by leveraging and fusing the complementary information from different time points. Our algorithm is highly flexible and allows training and inference on arbitrary number of low-resolution images. Experimental results show that our proposed SatDiffMoE method not only achieves superior performance for the satellite image super-resolution tasks on a variety of datasets, but also gets an improved computational efficiency with reduced model parameters, compared with previous methods.
@@ 1 Introduction
Satellite imaging is a very useful technique for monitoring the natural phenomena and human activities on the surfaces of the Earth. Lots of applications rely on satellite images such as crop monitoring, weather forecasting, urban planning, wildfire management and so on  [DiffusionSat, sustainabledevelopment, poverty1, poverty2, toefl, sat1, sat2, sat3]. However, the acquisition of satellite images can be very expensive and the spatial and temporal resolution (the frequency that a satellite image is captured) may be limited due to the physical constraints of sensors [DiffusionSat]. In addition, the high temporal resolution may come with trade-off in spatial resolution. Recent advance in satellite imaging technology enables us to capture the same area with a high-revisit frequency, but the spatial resolution is often limited. For instance, two Sentinel-2 satellites with resolutions from 10m to 60m can capture all land surfaces every five days [WorldStrat, satmae, sat4, sat5], but to perform land crop monitoring or wildfire management, this resolution is not sufficient. On the other hand, some very high-resolution satellite images such as SPOT6 or WorldView can have resolution better than 1.5m [fmow], but it is extremely difficult to collect those images for a large area, and these images cannot be captured as frequently as the Sentinel images. The limited temporal resolution severely limits downstream applications in urban planning, object detection, or continuous monitoring of crop or vegetation covers.
@@
To solve the aforementioned challenges, super-resolution algorithms have been introduced to predict the high-resolution (HR) satellite images from a bunch of corresponding low-resolution (LR) satellite images [fusion0, refusion1], so that to obtain more fine details. For example, given the low-resolution satellite images (10m) from Sentinel-2 of a specific location, the super-resolution algorithms are developed to predict the high-resolution (1.5m) SPOT6 satellite image of the same location at a specific time. Nevertheless, solving remote sensing super-resolution problems is still an open problem. One challenge in remote sensing is that low-resolution and high-resolution images often come from different sensors, and may maintain very different image features due to the large imaging modality gap resulted from different sensors [fusion0]. Moreover, since the low-resolution and high-resolution images are usually acquired at different time points as demonstrated in Fig. 1, a lot of atmospheric disturbance may pose additional challenges for modeling the sensor imaging process. Therefore, unlike the natural images, in the satellite images, the down-sampling process can be extremely difficult to model.
@@
In order to tackle this challenging task of remote sensing super-resolution, considering that the acquisition of satellite images often comes with multiple revisits (a collection of satellite images at the same location but at different time stamps) or with multiple spectral bands, we hypothesize that fusing multiple low-resolution images with different spectrums or at different time stamps may provide complementary information to the model so as to benefit the super-resolution task. With such motivation, existing works [WorldStrat] have introduced a recursive fusion module that takes the concatenated LR images as inputs and applies a residual attention model for outputting the HR image. DiffusionSat [DiffusionSat] introduces a 3D ControlNet architecture that fuses LR images of different spectral bands to reconstruct the corresponding HR image. However, these works require a fixed number of LR images or require an absolute timestamp for each LR and HR, which is often not feasible and flexible at inference time in real practice because it is challenging to find a fixed amount of paired LR images for each HR image.
@@ 2 Background
$dx_{t}=(-\frac{\beta(t)}{2}-\beta(t)\nabla_{x_{t}}\log p_{t}(x_{t}))dt+\sqrt{\beta(t)}d\overline{w}.$ (1)
@@
Latent diffusion models (LDM) [rombach2022highresolution] have been proposed for faster inference and training with a reduced computational burden. By applying an autoencoder to reduce data dimension, LDMs train the diffusion model in a compressed latent space, and then decode the latent code into signals. This method enables high-quality high-resolution image synthesis benefited from its compressed latent space, which is an ideal fit for satellite images due to the large image size. Nevertheless, it is still challenging to perform image restoration accurately with LDMs [ReSample, rombach2022highresolution]. Various works have tried to extend LDMs for high-dimensional or high-resolution signal synthesis, such as video generation [video1, video2, video3, video4, Video5, refusion2, refusion3, refusion4, refusion5]. However, few works apply LDMs for image fusion yet. It is an open problem to sample high-resolution images conditioning on multiple similar low-resolution images with LDMs.
@@
Single-image super-resolution (SISR) focuses on reconstructing the high-resolution image from one corresponding low-resolution image. In the era of deep learning, super-resolution problem has become a popular research question and many data-driven methods have been proposed. In past few years, state-of-the-art methods apply techniques such as Convolutional neural network (CNN), Generative Adversarial Network (GAN), and Transformers for restoring the high-resolution images [VAE, GAN, DBPN, MSRResNet, Pix2pix, Attentionallyouneed]. These methods usually learn the mapping from low-resolution image to high-resolution image through a data-driven way using the pair data to train the neural network.
However, many of these methods output blurry or inaccurate reconstructions since they are learning a direct mapping between LR images and HR images without considering the distribution of possible HR images [ReSample, PINER]. Due to the ill-posedness of the super-resolution problem, there may exist multiple HR images corresponding to one single LR image. A direct regression-based approach may let the network learn an average of all possible HR images, which leads to blurry output. Diffusion models address this issue by learning a strong generative prior that can perform posterior sampling [chung:23:dps] instead of direct regression. This sampling method outputs realistic images, and leads to better image perceptual quality. One line of work assumes the degradation operator is known, and focuses on inference-time posterior sampling with the diffusion prior without retraining [chung:23:dps, DDRM, DDNM, ReSample, Solve2]. The other line of work assume the degradation operator is unknown. They concat the LR image into the noise vector or conditional networks and then retrain or fine-tune the diffusion model [ControlNet, DiffusionSat]. Both lines of work show good ability to model complex high-resolution image distributions and outperform CNN approaches in image perceptual quality. However, all these works focus on obtaining HR images from a single LR image.
@@
The goal of multi-image super-resolution is to combine the information from multiple LR images to reconstruct one HR image. In satellite imagery, different sensors have different resolutions. For instance, Sentinel-2 SITS has a resolution of 10-60m, but SPOT-6 or fMoW can have a resolution of less than 1.5m [DiffusionSat]. Given sequential low-resolution images collected at the same location but different times, the hypothesis is that performing multi-image super-resolution is able to combine the information of LR images at multiple times to obtain a more accurate and higher-quality HR image. There are a couple of recent works in this venue [fusion0, WorldStrat, DiffusionSat]. For example, Cornebise et al. trained a network with a traditional autoencoder architecture, but modify the encoder to incorporate multiple images as input [WorldStrat]. HighRes-net [WorldStrat] adopted this idea to solve satellite image fusion problem with a network composed of an encoder, a recursive fusion network, and a decoder. However this approach does not include the temporal information of LR images DiffusionSat [DiffusionSat] proposed to train a 3D ControlNet on top of a fine-tuned latent diffusion model that leverages multispectral bands for reconstruction. However, at training, the 3D ControlNet requires the same number of low-resolution images for each paired high-resolution images. In addition, at inference-time, the number of low-resolution images used for reconstruction must be the same as the training set. Motivated by this, essentially different from all these previous works, we aim to propose a more flexible and robust fusion algorithm that can take an arbitrary number of low-resolution inputs with corresponding time information to generate the high-resolution satellite image.
@@ 3 Methods
$i\text{-}th$ low-resolution images $\text{LR}_{i}$, relative time difference $dt_{i}$, $i=1$, $\ldots$, N, Encoder $\mathcal{E}(\cdot)$, Decoder $\mathcal{D}(\cdot)$, Score function $\bm{s}_{\theta}(\cdot,t)$, Pretrained LDM parameters $\beta_{t}$, $\bar{\alpha}_{t}$, $\eta$, $\delta$, Hyperparameter $\lambda$ to control the fusion strength, $d(\cdot)$ the distance function.
$\bm{z}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$triangleright Initial noise vector
@@
     $\boldsymbol{\epsilon}_{1}\sim\mathcal{N}(\bm{0},\bm{I})$
     $\hat{\bm{\epsilon}}_{t+1}^{i}=\bm{s}_{\theta}(\bm{z}_{t+1}^{i},{t+1},\text{LR}_{i},dt_{i})$triangleright Compute the score
     $\hat{\bm{z}}_{0}^{i}(\bm{z}_{t+1}^{i})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}(\bm{z}_{t+1}^{i}-\sqrt{1-\bar{\alpha}_{t+1}^{i}}\hat{\bm{\epsilon}}_{t+1}^{i})$triangleright Predict $\hat{\bm{z}}_{0}$ using Tweedie’s formula
     $\bar{\bm{z}}_{0}\in\underset{\bm{z}}{\operatorname*{arg\,min}}\,\sum^{N}_{i=0}d(z,\hat{\bm{z}}_{0}^{i}(\bm{z}_{t+1}^{i}))$triangleright Find the center of multiple $\hat{\bm{z}}_{0}$ by optimization
     $\hat{\bm{z}}_{0}^{i}(\text{LR}_{i})=(1-\lambda)\hat{\bm{z}}_{0}^{i}(\bm{z}_{t+1}^{i}))+\lambda\bar{\bm{z}}_{0}$
     $\bm{z}_{t}^{i}=\sqrt{\bar{\alpha}_{t-1}}\hat{\bm{z}}^{i}_{0}(\text{LR}_{i})+\sqrt{1-\bar{\alpha}_{t-1}-\eta\delta^{2}}\hat{\bm{\epsilon}}_{t+1}^{i}+\eta\delta\boldsymbol{\epsilon}_{1}$triangleright Update intermediate noisy samples
$\bm{x}^{i}_{0}=\mathcal{D}(\bm{z}^{i}_{0})$triangleright Output reconstructed image
@@
Instead of conditioning on multiple low-resolution images as a concatenated input [DiffusionSat], we propose a novel fusion algorithm: condition on each image and then fuse each score (the output of the conditional diffusion model) into a high-resolution image reconstruction. First, in order to embed the temporal information, we introduce a new method that trains a conditional diffusion model using one LR image and the relative time difference (between the input LR image and the target HR image) as inputs to reconstruct the HR image. Then we propose a novel inference-time algorithm that modifies the intermediate outputs in the reverse sampling procedure based on the assumption that HR reconstruction of LR images at the same location should look similar. Our method is illustrated in Fig. 2.
@@
We propose a novel conditioning mechanism for reconstructing HR image from one LR image and its corresponding time difference with the target HR image. Firstly, we utilize Stable Diffusion [rombach2022highresolution], a pre-trained latent diffusion model for natural image generation, as the backbone model and fine tune the model on satellite images. We also propose a novel method to add additional control to the generating process.
@@
$\mathop{\arg\min}\limits_{\theta}\mathbb{E}_{\text{LR}_{i},\text{HR},t,\epsilon\in N(0,1)}||F(s_{\theta}((\text{concat}(\mathcal{E}(\text{LR}_{i}),\mathcal{E}(\text{HR}))_{t},t)-\epsilon)||^{2}_{2}$ (2)
@@
$\mathop{\arg\min}\limits_{\theta}\mathbb{E}_{\text{LR}_{i},\text{HR},t,\epsilon\in N(0,1)}||F(s_{\theta}((\text{concat}(\mathcal{E}(\text{LR}_{i}),\mathcal{E}(\text{HR}))_{t},dt_{i},t)-\epsilon)||^{2}_{2}$ (3)
@@
$\bar{z}_{0}=\mathop{\arg\min}\limits_{x}\sum_{i=1}^{N}d(x,\mathbb{E}[z_{0}|z_{t}(\text{LR}_{i},dt_{i})])$ (4)
@@
After obtaining $\bar{z}_{0}$, we propose to update all intermediate samples from the $\text{LR}_{i}$ to be closer to $\bar{z}_{0}$. Recall that in DDIM reverse sampling [DDIM], the reverse sampling can be decomposed by a clean image component and a noise component. We follow [DDS]’s approach that only updates the clean image component, and leaves the noise component intact. Specifically, let $\lambda$ be a hyperparameter balancing the original clean image component and $\bar{z}_{0}$, we can update the new clean component as:
@@
Therefore, the overall reverse sampling step can be written as
$\bm{z}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{z}_{0}(\text{LR}_{i},dt_{i})+\sqrt{1-\bar{\alpha}_{t-1}-\eta\delta_{t}^{2}}\bm{s}_{\theta}(\bm{z}_{t},\text{LR}_{i},dt_{i},t)+\eta\delta_{t}\bm{\epsilon},\quad t=T,\ldots,0,$ (7)

### v2

# SatDiffMoE: A Mixture of Estimation Method for Satellite Image Super-resolution with Latent Diffusion Models

@@ Abstract
During the acquisition of satellite images, there is generally a trade-off between spatial resolution and temporal resolution (acquisition frequency) due to the onboard sensors of satellite imaging systems. With the advent of diffusion models, we can now learn strong generative priors to generate realistic satellite images with high resolution, which can be utilized to promote the super-resolution task as well. In this work, we propose a novel diffusion-based fusion algorithm called SatDiffMoE that can take an arbitrary number of sequential low-resolution satellite images at the same location as inputs, and fuse them into one high-resolution reconstructed image with more fine details, by leveraging and fusing the complementary information from different time points. Our algorithm is highly flexible and allows training and inference on arbitrary number of low-resolution images. Experimental results show that our proposed SatDiffMoE method not only achieves superior performance for the satellite image super-resolution tasks on a variety of datasets, but also gets an improved computational efficiency with reduced model parameters, compared with previous methods.
Machine Learning, ICML
@@ 1 Introduction
Satellite imaging is a very useful technique for monitoring the natural phenomena and human activities on the surfaces of the Earth. Lots of applications rely on satellite images such as crop monitoring, weather forecasting, urban planning, wildfire management and so on  [Khanna et al., 2024, Burke et al., 2021, Ayush et al., 2020, 2021, Beck et al., 2007, Wang et al., 2018a, M Rustowicz et al., 2019, Li et al., 2019]. However, the acquisition of satellite images can be very expensive and the spatial and temporal resolution (the frequency that a satellite image is captured) may be limited due to the physical constraints of sensors [Khanna et al., 2024]. In addition, the high temporal resolution may come with trade-off in spatial resolution. Recent advance in satellite imaging technology enables us to capture the same area with a high-revisit frequency, but the spatial resolution is often limited. For instance, two Sentinel-2 satellites with resolutions from 10m to 60m can capture all land surfaces every five days [Cornebise et al., 2022, Cong et al., 2022, Tarasiewicz et al., 2023, Van Etten et al., 2018], but to perform land crop monitoring or wildfire management, this resolution is not sufficient. On the other hand, some very high-resolution satellite images such as SPOT6 or WorldView can have resolution better than 1.5m [Christie et al., 2018], but it is extremely difficult to collect those images for a large area, and these images cannot be captured as frequently as the Sentinel images. The limited temporal resolution severely limits downstream applications in urban planning, object detection, or continuous monitoring of crop or vegetation covers.
@@
To solve the aforementioned challenges, super-resolution algorithms have been introduced to predict the high-resolution (HR) satellite images from a bunch of corresponding low-resolution (LR) satellite images [Luo et al., 2018, 2023], so that to obtain more fine details. For example, given the low-resolution satellite images (10m) from Sentinel-2 of a specific location, the super-resolution algorithms are developed to predict the high-resolution (1.5m) SPOT6 satellite image of the same location at a specific time. Nevertheless, solving remote sensing super-resolution problems is still an open problem. One challenge in remote sensing is that low-resolution and high-resolution images often come from different sensors, and may maintain very different image features due to the large imaging modality gap resulted from different sensors [Luo et al., 2018]. Moreover, since the low-resolution and high-resolution images are usually acquired at different time points as demonstrated in Fig. 1, a lot of atmospheric disturbance may pose additional challenges for modeling the sensor imaging process. Therefore, unlike the natural images, in the satellite images, the down-sampling process can be extremely difficult to model.
@@
In order to tackle this challenging task of remote sensing super-resolution, considering that the acquisition of satellite images often comes with multiple revisits (a collection of satellite images at the same location but at different time stamps) or with multiple spectral bands, we hypothesize that fusing multiple low-resolution images with different spectrums or at different time stamps may provide complementary information to the model so as to benefit the super-resolution task. With such motivation, existing works [Cornebise et al., 2022] have introduced a recursive fusion module that takes the concatenated LR images as inputs and applies a residual attention model for outputting the HR image. DiffusionSat [Khanna et al., 2024] introduces a 3D ControlNet architecture that fuses LR images of different spectral bands to reconstruct the corresponding HR image. However, these works require a fixed number of LR images or require an absolute timestamp for each LR and HR, which is often not feasible and flexible at inference time in real practice because it is challenging to find a fixed amount of paired LR images for each HR image.
@@ 2 Background
$dx_{t}=\left(-\frac{\beta(t)}{2}-\beta(t)\nabla_{x_{t}}\log p_{t}(x_{t})\right)dt+\sqrt{\beta(t)}d\overline{w}.$ (1)
@@
Latent diffusion models (LDM) [Rombach et al., 2022] have been proposed for faster inference and training with a reduced computational burden. By applying an autoencoder to reduce data dimension, LDMs train the diffusion model in a compressed latent space, and then decode the latent code into signals. This method enables high-quality high-resolution image synthesis benefited from its compressed latent space, which is an ideal fit for satellite images due to the large image size. Nevertheless, it is still challenging to perform image restoration accurately with LDMs [Song et al., 2024, Rombach et al., 2022]. Various works have tried to extend LDMs for high-dimensional or high-resolution signal synthesis, such as video generation [Blattmann et al., 2023, Yu et al., 2023, Ni et al., 2023, Ho et al., 2022, Ceylan et al., 2023, Martinez et al., 2021, Ruiz et al., 2023, Saharia et al., 2022, Voleti et al., 2022]. However, few works apply LDMs for image fusion yet. It is an open problem to sample high-resolution images conditioning on multiple similar low-resolution images with LDMs.
@@
Single-image super-resolution (SISR) focuses on reconstructing the high-resolution image from one corresponding low-resolution image. In the era of deep learning, super-resolution problem has become a popular research question and many data-driven methods have been proposed. In past few years, state-of-the-art methods apply techniques such as Convolutional neural network (CNN), Generative Adversarial Network (GAN), and Transformers for restoring the high-resolution images [Kingma and Welling, 2013, Creswell et al., 2018, Haris et al., 2018, Wang et al., 2018b, Isola et al., 2017, Vaswani et al., 2017]. These methods usually learn the mapping from low-resolution image to high-resolution image through a data-driven way using the pair data to train the neural network.
However, many of these methods output blurry or inaccurate reconstructions since they are learning a direct mapping between LR images and HR images without considering the distribution of possible HR images [Song et al., 2024, 2023]. Due to the ill-posedness of the super-resolution problem, there may exist multiple HR images corresponding to one single LR image. A direct regression-based approach may let the network learn an average of all possible HR images, which leads to blurry output. Diffusion models address this issue by learning a strong generative prior that can perform posterior sampling [Chung et al., 2023] instead of direct regression. This sampling method outputs realistic images, and leads to better image perceptual quality. One line of work assumes the degradation operator is known, and focuses on inference-time posterior sampling with the diffusion prior without retraining [Chung et al., 2023, Kawar et al., 2022, Wang et al., 2022, Song et al., 2024, 2021]. The other line of work assume the degradation operator is unknown. They concat the LR image into the noise vector or conditional networks and then retrain or fine-tune the diffusion model [Zhang et al., 2023, Khanna et al., 2024]. Both lines of work show good ability to model complex high-resolution image distributions and outperform CNN approaches in image perceptual quality. However, all these works focus on obtaining HR images from a single LR image.
@@
The goal of multi-image super-resolution is to combine the information from multiple LR images to reconstruct one HR image. In satellite imagery, different sensors have different resolutions. For instance, Sentinel-2 SITS has a resolution of 10-60m, but SPOT-6 or fMoW can have a resolution of less than 1.5m [Khanna et al., 2024]. Given sequential low-resolution images collected at the same location but different times, the hypothesis is that performing multi-image super-resolution is able to combine the information of LR images at multiple times to obtain a more accurate and higher-quality HR image. There are a couple of recent works in this venue [Luo et al., 2018, Cornebise et al., 2022, Khanna et al., 2024]. For example, Cornebise et al. trained a network with a traditional autoencoder architecture, but modify the encoder to incorporate multiple images as input [Cornebise et al., 2022]. HighRes-net [Cornebise et al., 2022] adopted this idea to solve satellite image fusion problem with a network composed of an encoder, a recursive fusion network, and a decoder. However this approach does not include the temporal information of LR images DiffusionSat [Khanna et al., 2024] proposed to train a 3D ControlNet on top of a fine-tuned latent diffusion model that leverages multispectral bands for reconstruction. However, at training, the 3D ControlNet requires the same number of low-resolution images for each paired high-resolution images. In addition, at inference-time, the number of low-resolution images used for reconstruction must be the same as the training set. Motivated by this, essentially different from all these previous works, we aim to propose a more flexible and robust fusion algorithm that can take an arbitrary number of low-resolution inputs with corresponding time information to generate the high-resolution satellite image.
@@ 3 Methods
$i\text{-}th$ low-resolution images $\text{LR}_{i}$, relative time difference $dt_{i}$, $i=1$, $\ldots$, N, Encoder $\mathcal{E}(\cdot)$, Decoder $\mathcal{D}(\cdot)$, Score function $\mathbf{s}_{\theta}(\cdot,t)$, Pretrained LDM parameters $\beta_{t}$, $\bar{\alpha}_{t}$, $\eta$, $\delta$, Hyperparameter $\lambda$ to control the fusion strength, $d(\cdot)$ the distance function.
$\mathbf{z}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$triangleright Initial noise vector
@@
     $\boldsymbol{\epsilon}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
     $\hat{\mathbf{\epsilon}}_{t+1}^{i}=\mathbf{s}_{\theta}(\mathbf{z}_{t+1}^{i},{t+1},\text{LR}_{i},dt_{i})$triangleright Compute the score
     $\hat{\mathbf{z}}_{0}^{i}(\mathbf{z}_{t+1}^{i})=\frac{1}{\sqrt{\bar{\alpha}_{t+1}}}(\mathbf{z}_{t+1}^{i}-\sqrt{1-\bar{\alpha}_{t+1}^{i}}\hat{\mathbf{\epsilon}}_{t+1}^{i})$triangleright Predict $\hat{\mathbf{z}}_{0}$ using Tweedie’s formula
     $\bar{\mathbf{z}}_{0}\in\underset{\mathbf{z}}{\text{argmin}}\,\sum^{N}_{i=0}d(z,\hat{\mathbf{z}}_{0}^{i}(\mathbf{z}_{t+1}^{i}))$triangleright Find the center of multiple $\hat{\mathbf{z}}_{0}$ by optimization
     $\hat{\mathbf{z}}_{0}^{i}(\text{LR}_{i})=(1-\lambda)\hat{\mathbf{z}}_{0}^{i}(\mathbf{z}_{t+1}^{i}))+\lambda\bar{\mathbf{z}}_{0}$
     $\mathbf{z}_{t}^{i}=\sqrt{\bar{\alpha}_{t-1}}\hat{\mathbf{z}}^{i}_{0}(\text{LR}_{i})+\sqrt{1-\bar{\alpha}_{t-1}-\eta\delta^{2}}\hat{\mathbf{\epsilon}}_{t+1}^{i}+\eta\delta\boldsymbol{\epsilon}_{1}$triangleright Update intermediate noisy samples
end for
$\mathbf{x}^{i}_{0}=\mathcal{D}(\mathbf{z}^{i}_{0})$triangleright Output reconstructed image
@@
Instead of conditioning on multiple low-resolution images as a concatenated input [Khanna et al., 2024], we propose a novel fusion algorithm: condition on each image and then fuse each score (the output of the conditional diffusion model) into a high-resolution image reconstruction. First, in order to embed the temporal information, we introduce a new method that trains a conditional diffusion model using one LR image and the relative time difference (between the input LR image and the target HR image) as inputs to reconstruct the HR image. Then we propose a novel inference-time algorithm that modifies the intermediate outputs in the reverse sampling procedure based on the assumption that HR reconstruction of LR images at the same location should look similar. Our method is illustrated in Fig. 2.
@@
We propose a novel conditioning mechanism for reconstructing HR image from one LR image and its corresponding time difference with the target HR image. Firstly, we utilize Stable Diffusion [Rombach et al., 2022], a pre-trained latent diffusion model for natural image generation, as the backbone model and fine tune the model on satellite images. We also propose a novel method to add additional control to the generating process.
@@
$\mathop{\arg\min}\limits_{\theta}\mathbb{E}_{\text{LR}_{i},\text{HR},t,\epsilon\in N(0,1)}||F(s_{\theta}$
$((\text{concat}(\mathcal{E}(\text{LR}_{i}),\mathcal{E}(\text{HR}))_{t},t)-\epsilon)||^{2}_{2}$ (2)
@@
$\mathop{\arg\min}\limits_{\theta}\mathbb{E}_{\text{LR}_{i},\text{HR},t,\epsilon\in N(0,1)}||F(s_{\theta}$
$((\text{concat}(\mathcal{E}(\text{LR}_{i}),\mathcal{E}(\text{HR}))_{t},dt_{i},t)-\epsilon)||^{2}_{2}$ (3)
@@
$\bar{z}_{0}=\mathop{\arg\min}\limits_{x}\sum_{\mathclap{i=1}}^{N}d(x,\mathbb{E}[z_{0}|z_{t}(\text{LR}_{i},dt_{i})])$ (4)
@@
After obtaining $\bar{z}_{0}$, we propose to update all intermediate samples from the $\text{LR}_{i}$ to be closer to $\bar{z}_{0}$. Recall that in DDIM reverse sampling [Song et al., 2022], the reverse sampling can be decomposed by a clean image component and a noise component. We follow [Chung et al., 2024]’s approach that only updates the clean image component, and leaves the noise component intact. Specifically, let $\lambda$ be a hyperparameter balancing the original clean image component and $\bar{z}_{0}$, we can update the new clean component as: